xx IDWaveMoL

IDWaveMoL

Ian Hawke

Date

Abstract

Initial data for the CactusExamples/WaveMoL thorn.

1 Using This Thorn IDWaveMoL provides simple initial data to the WaveMoL thorn. Currently only a time symmetric gaussian pulse is provided.

2 Parameters

11 amplitude 10 Scope: private01  REAL

11 Description: The amplitude of the waves

11 Range 10              01  Default: 1.0
11 *:*  10 No restriction

11

11 centrex 10 Scope: private01  REAL

11 Description: The centre of the gaussian wave, x-direction

11 Range 10              01  Default: 0.0
11 *:*  10 Anything

11

11 centrey 10 Scope: private01  REAL

11 Description: The centre of the gaussian wave, y-direction

11 Range 10              01  Default: 0.0
11 *:*  10 Anything

11

11 centrez  10 Scope: private01  REAL

11 Description: The centre of the gaussian wave, z-direction

11 Range 10              01  Default: 0.0
11 *:*  10 Anything

11

11 kx  10 Scope: private01  REAL

11 Description: Wave number in the x-direction

11 Range 10              01  Default: 0.0
11 *:*  10 Anything

11

11 ky  10 Scope: private01  REAL

11 Description: Wave number in the y-direction

11 Range 10              01  Default: 0.0
11 *:*  10 Anything

11

11 kz  10 Scope: private01  REAL

11 Description: Wave number in the z-direction

11 Range 10              01  Default: 0.0
11 *:*  10 Anything

11

11 offsett  10 Scope: private01  REAL

11 Description: Wave offset in the time-direction

11 Range 10              01  Default: 0.0
11 *:*  10 Anything

11

11 offsetx  10 Scope: private01  REAL

11 Description: Wave offset in the x-direction

11 Range 10              01  Default: 0.0
11 *:*  10 Anything

11 offsety  10 Scope: private01  REAL

11 Description: Wave offset in the y-direction

11 Range 10              01  Default: 0.0
11 *:*  10 Anything

11

11 offsetz  10 Scope: private01  REAL

11 Description: Wave offset in the z-direction

11 Range 10              01  Default: 0.0
11 *:*  10 Anything

11

11 radius  10 Scope: private01  REAL

11 Description: The radius of the gaussian wave

11 Range 10              01  Default: 0.0
11 0:*  10 Positive

11

11 sigma  10 Scope: private01  REAL

11 Description: The sigma for the gaussian wave

11 Range 10              01  Default: 0.1
11 0:*  10 Positive

11

11 slopet  10 Scope: private01  REAL

11 Description: Wave slope in the time-direction

11 Range 10              01  Default: 0.0
11 *:*  10 Anything

11

11 slopex  10 Scope: private01  REAL

11 Description: Wave slope in the x-direction

11 Range 10              01  Default: 0.0
11 *:*  10 Anything

11

11 slopey  10 Scope: private01  REAL

11 Description: Wave slope in the y-direction

11 Range 10              01  Default: 0.0
11 *:*  10 Anything

11

11 slopez  10 Scope: private01  REAL

11 Description: Wave slope in the z-direction

11 Range 10              01  Default: 0.0
11 *:*  10 Anything

11 initialxdata10 Scope: restricted01  KEYWORD

11 Description: Type of initial data

11 Range 10                01  Default: gaussian
11 gaussian 10 Gaussian wave
11 plane  10 plane wave

11

3 Interfaces

General Implements: idwavemol

Inherits: wavemol grid

4 Schedule This section lists all the variables which are assigned storage by thorn CactusExamples/IDWaveMoL. Storage can either last for the duration of the run (Always means that if this thorn is activated storage will be assigned, Conditional means that if this thorn is activated storage will be assigned for the duration of the run if some condition is met), or can be turned on for the duration of a schedule function.

Storage NONE

Scheduled Functions CCTKxPARAMCHECK x idwavemolxcheckparameters x check parameters x

11 x 11 Language: 11 c
11 x 11 Type:  11 function

11

CCTKxINITIAL x idwavemolxinitialdata x initial data for 3d wave equation x

11 x 11 Language: 11 c
11 x 11 Reads:  11 grid::x
11 x 11 x  11 grid::y
11 x 11 x  11 grid::z
11 x 11 Storage: 11 wavemol::scalarevolvemolxscalar[2]
11 x 11 x  11 wavemol::scalarevolvemolxvector[2]
11 x 11 Type:  11 function
11 x 11 Writes: 11 wavemol::scalarevolvemolxscalar(everywhere)
11 x 11 x  11 wavemol::scalarevolvemolxvector(everywhere)